2 cos 30 degrees + cot 60 degrees

2 sin 45 degrees - csc 45 degrees


1
Expert's answer
2021-06-15T16:03:37-0400

"a)\\
2cos30°+cot60°\\
=2cos30°+\frac{cos60°}{sin60°}\\
=2\frac{\sqrt3}{2}+\frac{\frac{1}{2}
}{\frac{\sqrt3}{2}}\\
=\sqrt3+\frac{1}{\sqrt3}\\
b)\\
2sin45°-csc45°\\
=2sin45°-\frac{1}{sin45°}\\
=2×\frac{1}{\sqrt2}-\frac{1}{\frac{1}{\sqrt3}}\\
=\sqrt2-\sqrt3"